You are a financial investor who buys and sells in the securities market. Now you have a portfolio of all blue chips including $11600 of share A, $7800 of share B, $14900 of share C and $ 3200 of share D, respectively. (A) Compute the weight of the assets in your portfolio. (B) If your portfolio has provided you with returns 7.6%, 12.2%, -4.7% and 13.4% over the past four years respectively. Calculate the geometric average return of the portfolio for this period. (C) Assume that expected return of the stock A in your portfolio is 15.2%. The risk premiums of the stock of the same industry are 4.8%, betas of these stocks is 1.3 and the inflation rate was 4.7% calculate the risk free rate of return using capital market pricing model (CAPM) You have another portfolio that comprises of two shares only 1200 golden sand shares and 400 silver beach shares. Below is the data of the portfolio. Expect return 12%. 17% Correlation of coefficient (p) 20% 40% Compute the expected return of your portfolio (E) Compute the expected risk 9standard deviation) of portfolio. (F) You bought those two shares two years ago with total investment of $16000.golden sand paid a dividend of $4.4 / share per year. Silver beach paid a dividend of $7.5/share per year. Calculate your total capital gain of this portfolio if you today can sell golden sand for $12 per share and silver black for $18 per share.
Answer #1

Let me tell you upfront that part (D) & (E) can't be answered. I have answered all other parts. What's required to solve has been mentioned at an appropriate place. Please go through the solution. Please don't give a negative rating merely because I believe there is insufficient information to answer part (D) and (E)

Part (A)

Stock Amount Proportion Calculated as
A             11,600 30.93% = 11600 / 37500
B                7,800 20.80% = 7800 / 37500
C             14,900 39.73% = 14900 / 37500
D                3,200 8.53% = 3200 / 37500
Total             37,500

Part (B)

If your portfolio has provided you with returns 7.6%, 12.2%, -4.7% and 13.4% over the past four years respectively, then the geometric average return of the portfolio for this period = [(1 + 7.6%) x (1 + 12.2%) x (1 - 4.7%) x (1 + 13.4%)]1/4 -1 = 6.88%

Part (C)

Risk free rate = Expected return - Beta x market premium = 15.2% - 1.3 x 4.8% = 8.96%

Part (D) can't be answered as there is ambiguity. The number of shares have been given. But the price per share or portfolio weight is not available. In the absence of these information weights of the two shares can't be calculated and hence the answer can't be calculated.

Part (E) can't be answered because there is an ambiguity in the question. Correlation of coefficient is given but there is no information on standard deviation. Further, there has to be one figure of coefficient of correlation between the two stocks. It seems some information has been lost. Please check if these two sub parts are complete in all respect.

Part (F)

Value of the portfolio now = 12 x 1,200 + 18 x 400 = $ 21,600

Hence, total capital gain = 21,600 - 16,000 = $ 5,600